# Efficient Computation of Closed Substrings

### v1

@@ Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present a fast and practical $\mathcal{O}(n\log n)$ time algorithm to compute all $\Theta(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. We also present a simple and space-efficient solution to compute all maximal closed substrings (MCSs) using the suffix array ($\mathsf{SA}$) and the longest common prefix ($\mathsf{LCP}$) array of $w[1..n]$. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
losed Strings, Maximal Closed Substrings, and Fibonacci Words
Keywords:
C
@@ 1 Introduction
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$; otherwise, it is called open. In other words, it is the longest border of $u$ that does not have internal occurrences in $u$. For example, in the string $abcab$, $ab$ occurs only as the prefix and suffix. Note that this longest border may overlap. A substring $u$ of a string $w$ is said to be maximal right (left)-closed if it is a closed string and cannot be extended to the right (left) by a character to form another closed string. A maximal closed substring (MCS) of a string $w$ is a closed string that is both maximal left and maximal right-closed. An MCS of length one is called a singleton MCS; otherwise, it is called a non-singleton MCS. For example, given a string $w=abaccaba$, the non-singleton MCSs are $abaccaba$, $acca$, $aba$, $cc$, and the singleton MCSs are $a$ and $b$. Observe that $baccab$, $baccaba$, $abaccab$, and $c$ are closed but not maximal.
@@
It has been shown that a string contains at most $\Theta(n^{2})$ distinct closed factors [5]. More recently (in 2024), [19] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [16] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\Theta(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using sliding suffix trees and weighted ancestor queries factors.
In [4] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [6] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later proved by Kosolobov in [14].
@@
Compact Representation of All $\Theta(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we define a compact representation of all $\Theta(n^{2})$ closed substrings of $w$ using $\mathcal{O}(n\log n)$ space.
Compute All $\Theta(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we find all $\Theta(n^{2})$ closed substrings of $w$ using suffix arrays ($\mathsf{SA}$) and longest common prefix arrays ($\mathsf{LCP}$) in $\mathcal{O}(n\log n)$ time.
@@
Exact Number of MCSs in a Fibonacci Word: Given a Fibonacci word $f_{n}$, the number of MCSs in $f_{n}$ is denoted by $M(f_{n})$ and is defined in Equation (8). We also show that $M(f_{n})\approx 1.382F_{n}$.
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present algorithms to efficiently compute all closed and MCSs in $\mathcal{O}(n\log n)$ time. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. Finally, implementation details and experiments are discussed in Section 6.
@@ 2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n=|w|$. An empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if $j-i+1<n$. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover. If a proper cover $u$ exists, then we say that $w$ is quasiperiodic (for more details, see the survey [15]).
A string of the form $u^{e}$, where $u\neq\varepsilon$, $e\in\mathbb{Z}$, and $e\geq 2$ is called a repetition; otherwise, it is called primitive. A string $w[1..n]$ is said to be periodic if it can be represented as $w=u^{e}u^{\prime}$, where $u^{\prime}$ is a (possibly empty) prefix of $u$, and either $e\geq 2$ or $e=1$ with $u^{\prime}$ non-empty. Here, $e$ is called exponent, $p=|u|$ is called the period, and $u$ is called the generator of $w$. A substring $u=w[i..j]$ is called a run if it satisfies two conditions: first, $u=v^{e}v^{\prime}$, where $e\geq 2$, $v$ is primitive, and $v^{\prime}$ is a (possibly empty) prefix of $v$; and second, the periodicity cannot be extended to the left or right, that is, $i=1$ or $w[i-1]\neq w[i+p-1]$, and $j=n$ or $w[j+1]\neq w[j-p+1]$. Here, $v$ is called the base of the run.
A gapped repeat is a substring of $w[1..n]$ of the form $uvu$, where $u$ and $v$ are non-empty strings. The substring $u$ is called the arm of the gapped repeat. The period of the gapped repeat is $|u|+|v|$. The gapped repeat is called an $\alpha$-gapped if its period is not greater than $\alpha|u|$, where $\alpha\geq 1$. An $\alpha$-gapped repeat is called a maximal $\alpha$-gapped if it cannot be extended to the left or to the right by at least one character while preserving its period. A gapped-MCS is a maximal $\alpha$-gapped repeat whose arm occurs exactly twice.
@@
Given a string $w[1..n]$, we say that a closed substring $u$ of $w$ starting at index $i$ is right-extendible to a maximal right-closed substring $r$ of $w$, if there exists a $u^{\prime}$ such that $r=uu^{\prime}$, where $u^{\prime}$ could be $\varepsilon$. We denote the maximal right-closed substring, the border, and the shortest closed prefix by $r$, $b$, and $p$, respectively, and their lengths by $\ell r$, $\ell b$, and $\ell p$, respectively.
@@ 3 Compact Representation of All Closed Substrings
A string $w[1..n]$ contains $\Theta(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\Theta(n^{2})$ closed substrings in Theorem 3.1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 2 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$. Fact 1 follows from the definitions of closed and maximal right-closed substring.
Fact 1
A closed substring $u=w[i..j]$ of a string $w[1..n]$ is either maximal right-closed or there exists a unique maximal right-closed substring $r=w[i..j^{\prime}]$ such that $u$ is right-extendible to $r$. Moreover, every prefix $v=w[i..k]$ of $r$ is also closed, where $1\leq i\leq j\leq j^{\prime}\leq n$ and $j<k\leq j^{\prime}$.
By Fact 1, since closed prefixes are right-extendible to a unique maximal right-closed substring, in Equation (1) we define the equivalence class $E_{r}$ of a maximal right-closed substring $r$ of $w[1..n]$, where $u$ and $v$ are closed prefixes of $r$.
$(u,v)\in E_{r}\iff u\text{ and }v\text{ are right-extendible to }r$ (1)
@@
For $1\leq m\leq n$, $w[1..m]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(m=n\land w[1..m]=\lambda^{m})\lor(m<n\land w[1..m+1]=\lambda^{m}\lambda_{1})$, where $\lambda,\lambda_{1}\in\Sigma$, and $\lambda\neq\lambda_{1}$. Moreover, each prefix of $w[1..m]$ is closed.
@@
We prove the lemma in two parts. First, we show by induction on $m$ that for $1\leq m\leq n$, $w[1..m]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(m=n\land w[1..m]=\lambda^{m})\lor(m<n\land w[1..m+1]=\lambda^{m}\lambda_{1})$.
The base case, where $m=1$, trivially holds. For the inductive step, assume that for some $k\geq 2$, $w[1..k]$ is the shortest maximal right-closed prefix of string $w[1..n]$ $\iff$ $(k=n\land w[1..k]=\lambda^{k})\lor(k<n\land w[1..k+1]=\lambda^{k}\lambda_{1})$.
()($\implies$)( ⟹ ) If $k+1<n$, suppose $w[1..k+1]$ is the shortest maximal right-closed prefix of $w[1..n]$. Assume for contradiction that $w[1..(k+1)+1]\neq\lambda^{k+1}\lambda_{1}$. Let $j$ be the shortest index such that $w[j]=\lambda_{2}$, where $\lambda_{2}\neq\lambda_{1}\neq\lambda\in\Sigma$. Then, by the inductive hypothesis, the prefix $w[1..j-1]=\lambda^{j-1}$ is a maximal right-closed prefix of $w$. However, since $j-1<k+1$, $w[1..j-1]$, is the shortest maximal right closed prefix of $w$—a contradiction. Therefore, we must have that $w[1..k+1]=\lambda^{k+1}\lambda_{1}$.
If $k+1=n$. Let $w[1]=\lambda$. Then by definition of the shortest maximal right closed prefix, we have $w[1..k+1]=\lambda^{k+1}$.
@@
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to $j-1$, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed. ∎
Let $r_{1},r_{2},\ldots,r_{K_{i}}$ be all the maximal right-closed substrings starting at index $i$ of a string $w[1..n]$, where $1\leq i\leq n$, and $K_{i}$ is the number of maximal right-closed substrings starting at index $i$. Let the lengths of $r_{1},r_{2},\ldots,r_{K_{i}}$ be $\ell r_{1},\ell r_{2},\ldots,\ell r_{K_{i}}$, respectively, and ordered such that $\ell r_{1}<\ell r_{2}<\cdots<\ell r_{K_{i}}$.
In Equation (2) we give a formula to compute the length, $\ell p_{j}$, of the shortest closed prefix $p_{j}$ of $r_{j}$ such that $p_{j}\in E_{r_{j}}$, where $1\leq j\leq K_{i}$.
$\ell p_{j}=\begin{cases}1,&\text{if }j=1,\\ \ell r_{j}-\ell b_{j}+\ell b_{j-1}+1,&\text{if }1<j\leq K_{i}.\end{cases}$ (2)
where $b_{j}$ and $b_{j-1}$ denote the longest borders of $r_{j}$ and $r_{j-1}$, respectively, and their lengths are given by $\ell b_{j}$ and $\ell b_{j-1}$. Lemma 2 proves Equation (2) and is illustrated in Figure 1.
$r_{j}=p_{j}p^{\prime}_{j}$ of length $\ell r_{j}$$\ell b_{j-1}$$\ell b_{j-1}$$\ell b_{j-1}$$\ell b_{j}$$r_{j-1}$ of length $\ell r_{j-1}$$\ell b_{j-1}+1$$p_{j}$ of length $\ell p_{j}$$p^{\prime}_{j}$ of length $\ell p^{\prime}_{j}$$i$......$b_{j-1}$$b_{j-1}$$b_{j-1}$
Figure 1: Illustration of the second case in the proof of Lemma 2, showing that the length $\ell p_{j}$ of the shortest closed prefix $p_{j}$ of $r_{j}$ can be computed from the border lengths $\ell b_{j}$, $\ell b_{j-1}$ and the length of $r_{j}$ equal to $\ell r_{j}$, such that $p_{j}$ is right-extendible to $r_{j}=p_{j}p^{\prime}_{j}$, where $p^{\prime}_{j}$ could possibly be $\varepsilon$.
@@
The prefixes of $r_{j}$ with lengths in the interval $[\ell p_{j},\ell r_{j}]$ constitute the equivalence class $E_{r_{j}}$.
@@
We prove the Lemma by considering two cases:
For $j=1$, by Fact 1 and Lemma 1, all prefixes of $r_{1}$ are closed and constitute the equivalence class $E_{r_{1}}$.
For $1<j\leq K_{i}$, by definition, the shortest closed prefix $p_{j}$ starting at index $i$ and longer than $r_{j-1}$ (whose longest border has length $\ell b_{j-1}$) must have the longest border of length exactly $\ell b_{j-1}+1$. Let $|p_{j}|=\ell p_{j}$. By Fact 1, $p_{j}$ extends to $r_{j}=p_{j}p_{j}^{\prime}$ with the longest border $b_{j}$, and each prefix of $r_{j}$ with length at least $\ell p_{j}$ is closed. Since the length of the extension $p_{j}^{\prime}$ is $\ell p_{j}^{\prime}=\ell b_{j}-(\ell b_{j-1}+1)$, we have $\ell r_{j}=\ell p_{j}+\ell p_{j}^{\prime}=\ell p_{j}+\ell b_{j}-(\ell b_{j-1}+1)$. Solving for $\ell p_{j}$, we obtain the desired formula: $\ell p_{j}=\ell r_{j}-\ell b_{j}+\ell b_{j-1}+1$. ∎
@@
Theorem 3.1 follows directly from Fact 1, Lemma 2 and 3.
@@
[…]
where $K_{i}$ is the number of maximal right-closed substrings starting at position $i$, and for each $j$-th such substring $r_{j}$ of length $\ell r_{j}$, starting at index $i$, $\ell p_{j}$ denotes the length of its shortest closed prefix that can be right-extended to $r_{j}$.
@@
$\text{Closed}(w)=\bigcup_{(i,\ell p_{j},\ell r_{j})\in\mathcal{C}(w)}\left\{w[i\ldots i+\ell-1]\mid\ell\in[\ell p_{j},\ell r_{j}]\right\},$ (4)

### v2

@@ Abstract
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$. In this paper, we present a fast $\mathcal{O}(n\log n)$ time algorithm to compute all $\mathcal{O}(n^{2})$ closed substrings by introducing a compact representation for all closed substrings of a string $w[1..n]$, using only $\mathcal{O}(n\log n)$ space. We also present a simple and space-efficient solution to compute all maximal closed substrings (MCSs) using the suffix array ($\mathsf{SA}$) and the longest common prefix ($\mathsf{LCP}$) array of $w[1..n]$. Finally, we show that the exact number of MCSs ($M(f_{n})$) in a Fibonacci word $f_{n}$, for $n\geq 5$, is $\approx\left(1+\frac{1}{\phi^{2}}\right)F_{n}\approx 1.382F_{n}$, where $\phi$ is the golden ratio.
@@ 1 Introduction
A closed string $u$ is either of length one or contains a border that occurs only as a prefix and as a suffix in $u$ and nowhere else within $u$; otherwise, it is called open. In other words, the longest (possibly overlapping) border of a closed substring $u$ does not have internal occurrences in $u$. For example, in the string abcababcab, abab occurs only as a prefix and a suffix. An occurrence of a substring $u$ of a string $w$ is said to be maximal right (left)-closed if it is a closed string and cannot be extended to the right (left) by a character to form another closed string. A maximal closed substring (MCS) of a string $w$ is a substring that is both maximal left and maximal right-closed. An MCS of length one is called a singleton MCS; otherwise, it is called a non-singleton MCS. For example, in string $w[1..8]=abaccaba$, the non-singleton MCSs are $w[1..8]=abaccaba$, $w[3..6]=acca$, $w[1..3]=w[6..8]=aba$, $w[4..5]=cc$, and the singleton MCSs are all occurrences of $a$ and $b$. Observe that $w[2..7]=baccab$, $w[2..8]=baccaba$, $w[1..7]=abaccab$, and both occurrences of $c$ are closed but not maximal.
@@
It has been shown that a string contains at most $\mathcal{O}(n^{2})$ distinct closed factors [5]. More recently (in 2024), [23] showed that the maximal number of distinct closed factors in a word of length $n$ is $\frac{n^{2}}{6}$. In early 2025, [20] presented a linear algorithm to count the number of distinct closed factors of a string in $\mathcal{O}(n\log\sigma)$ time and another in $\mathcal{O}(n)$ time. They also described methods to enumerate all closed and distinct closed factors, achieving a complexity of $\mathcal{O}(n^{2})$, which is further optimized to $\mathcal{O}(n\sqrt{\log n}+\text{output}\cdot\log n)$ for distinct closed substrings, using weighted ancestor queries.
In [4] Badkobeh et al. introduced the notion of MCSs and proposed a $\mathcal{O}(n\sqrt{n})$ loose upper bound on the number of MCSs in a string of length $n$. They also proposed an algorithm using suffix trees to compute the MCSs for strings on a binary alphabet in $\mathcal{O}(n\log n+n\sqrt{n})$ time. Badkobeh et al. in [6] extended their algorithm to a general alphabet with a time complexity of $\mathcal{O}(n\log n)$, implicitly improving the bound on the total number of MCSs, this was later directly proved by Kosolobov in [18].
@@
Compact Representation of All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$ we define a compact representation of all $\mathcal{O}(n^{2})$ closed substrings of $w$ using $\mathcal{O}(n\log n)$ space.
Compute All $\mathcal{O}(n^{2})$ Closed Substrings: Given a string $w[1..n]$, we compute all $\mathcal{O}(n^{2})$ closed substrings of $w$ in a compact representation using the suffix array ($\mathsf{SA}$) and the longest common prefix array ($\mathsf{LCP}$), in $\mathcal{O}(n\log n)$ time.
@@
Exact Number of MCSs in a Fibonacci Word: Given a Fibonacci word $f_{n}$, we give an exact equation to find the number of MCSs in $f_{n}$ is denoted by $M(f_{n})$ and is defined in Equation (8). We also show that $M(f_{n})\approx 1.382F_{n}$.
In Section 2, we introduce the terminology and data structures used in the paper. In Section 3 we define a compact representation of all closed substrings of a given string $w[1..n]$. In Section 4, we present algorithms to efficiently compute all closed substrings in a compact representation, as well as all MCSs, in $\mathcal{O}(n\log n)$ time. In Section 5 we provide a formula to compute the exact number of MCSs in a Fibonacci word. Finally, we present experimental analysis in Section 6.
@@ 2 Preliminaries
An alphabet $\Sigma$, of size $\sigma$, is a set of symbols that is totally ordered. A string (word), written as $w[1..n]$, is an element of $\Sigma^{*}$, whose length is represented by $n$. The empty string of length zero is denoted by $\varepsilon$. A substring (factor) $w[i..j]$ of $w$ is a string, where $1\leq i,j\leq n$, if $j>i$ then the substring is $\varepsilon$. The substring $w[i..j]$ is called a proper substring, if ji+1<nj-i+1<n. A substring $w[i..j]$ where $i=1$ ($j=n$) is called a prefix (suffix) of $w$. If a substring $u$ of $w$ is both a proper prefix and a proper suffix of $w$, then it is called a border. A cover of a string $w$ is a substring $u$ of $w$ such that $w$ can be constructed by overlapping and/or adjacent instances of $u$. If $|u|<|w|$, then it is said to be a proper cover (for more details on covers, see the survey [19]).
A string $w[1..n]$ is periodic with period $t^{\prime}$ if $w[i]=w[i+t^{\prime}]$ holds for all $1\leq i\leq n-t^{\prime}$. The shortest period of $w$ is the smallest positive integer $t\leq t^{\prime}$ for which this condition holds. A periodic substring $w[i..j]$ of string $w[1..n]$ with shortest period $t$ is called a run if its length $j-i+1\geq 2t$ and its periodicity cannot be extended to the left or right, that is, $i=1$ or $w[i-1]\neq w[i+t-1]$, and $j=n$ or $w[j+1]\neq w[j-t+1]$.
A gapped repeat is a substring of $w[1..n]$ of the form uvuuvu, where $u$ and $v$ are non-empty strings. The substring $u$ is called the arm of the gapped repeat. The period [^1] Note that the term period, as used in the definition of a gapped repeat, differs from its earlier usage in the context of periodic strings. Throughout this paper, the term period is consistently used in reference to periodic strings or runs. of the gapped repeat is $|u|+|v|$. An occurrence of a gapped repeat is called a maximal gapped repeat if it cannot be extended to the left or to the right by at least one symbol while preserving its period. A maximal gapped repeat, which is also closed, is referred to as a gapped-MCS.
@@
A closed substring $u=w[i..j]$ of a string $w[1..n]$ is said to be right-extendible to a maximal right-closed substring $r=w[i..j^{\prime}]$ if $j<j^{\prime}\leq n$ and every prefix $w[i..k]$ of $r$ with $j<k<j^{\prime}$ is also closed. In this case, $r$ is the unique maximal right-closed extension of $u$.
@@ 3 Compact Representation of all Closed Substrings
A string $w[1..n]$ contains $\mathcal{O}(n^{2})$ closed substrings, which can naively be computed in $\mathcal{O}(n^{2})$ time by computing the border of all substrings of $w$. In this section, we define a compact representation for all $\mathcal{O}(n^{2})$ closed substrings in Theorem 3.1 that requires only $\mathcal{O}(n\log n)$ space. Algorithm 2 in Section 4 presents an $\mathcal{O}(n\log n)$ time solution to compute all closed substrings of $w$ in a compact representation of size $\mathcal{O}(n\log n)$.
Let $u$ and $v$ be the proper closed prefixes of the maximal right-closed substring $r$. Then, by definition of right-extendibility we define an equivalence relation $R$ as follows:
$R=\{(u,v)\mid u\text{ and }v\text{ are both right-extendible to }r\}.$
Next, we define the equivalence class $E_{r}$ of $r$ under the relation $R$ as follows:
$E_{r}=\{u\mid u\text{ is right-extendible to }r\}.$ (1)
@@
Given a string $w[1..n]$, if $w[1..m]$ is the shortest maximal right-closed prefix of $w$, then $w[1..m]=\lambda^{m}$, where $\lambda\in\Sigma$ and $1\leq m\leq n$. Moreover, each prefix of $\lambda^{m}$ is also closed.
@@
We prove the lemma in two parts. First part by contradiction. Suppose that $w[1..m]$ is the shortest maximal right-closed prefix of $w$, and that $w[1..m]\neq\lambda^{m}$. W.l.o.g let $w[1]=\lambda$. Let $1<j\leq m$ be the smallest index such that $w[j]=\lambda_{1}$ and $w[1..j-1]=\lambda^{j-1}$, where $\lambda_{1}\neq\lambda\in\Sigma$. Then, $w[1..j]=\lambda^{j}\lambda_{1}$. However, by definition of maximal right-closed string, $w[1..j-1]=\lambda^{j-1}$ is a shorter maximal right-closed string of $w$ — a contradiction.
Second, for any string of the form $u=\lambda^{m}$, consider its prefixes $p=u[1..j]$, where $1\leq j\leq m$. If $j=1$, then $p$ is closed trivially. If $j\geq 2$, then the border length of each prefix $p$ is equal to j1j-1, which is maximum and unique. Thus, all the prefixes of $u=\lambda^{m}$ are closed. ∎
Let $r$ denote a maximal right-closed substring, $b$ its border, and $p$ its shortest closed prefix such that $p=r$ or $p\in E_{r}$. Let $r_{1},r_{2},\ldots,r_{K_{i}}$ be all the maximal right-closed substrings starting at index $i$ of a string $w[1..n]$, where $1\leq i\leq n$, and $K_{i}$ is the number of maximal right-closed substrings starting at index $i$, such that $|r_{1}|<|r_{2}|<\cdots<|r_{K_{i}}|$.
@@
Figure 1: Illustration of the second case in the proof of Lemma 2, showing that $|p_{j}|$ can be computed from $|b_{j}|$, $|b_{j-1}|$ and $|r_{j}|$, such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$. Note that when $p_{j}=r_{j}$, $p_{j}^{\prime}=\varepsilon$.
In Equation (2) we present a formula to compute the length of the shortest closed prefix $p_{j}$ of $r_{j}$ such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$, where $1\leq j\leq K_{i}$. Lemma 2 proves Equation (2) and is illustrated in Figure 1.
@@
Let $1\leq j\leq K_{i}$. The length of the shortest closed prefix $p_{j}$ of $r_{j}$, such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$, is given by:
$|p_{j}|=\begin{cases}1,&\text{if }j=1,\\ |r_{j}|-|b_{j}|+|b_{j-1}|+1,&\text{if }1<j\leq K_{i}.\end{cases}$ (2)
where $b_{j}$ and $b_{j-1}$ are the longest borders of $r_{j}$ and $r_{j-1}$, respectively.
@@
We prove the Lemma by considering two cases. For $j=1$, by Equation (1) and Lemma 1, all prefixes of $r_{1}$ are closed and $|p_{1}|=1$.
For $1<j\leq K_{i}$, by definition of a closed substring, the shortest closed prefix $p_{j}$ starting at index $i$ and longer than $r_{j-1}$ (whose longest border has length $|b_{j-1}|$) must have the longest border of length exactly $|b_{j-1}|+1$. Either $p_{j}=r_{j}$ or by definition of right-extendibility $p_{j}$ extends to $r_{j}=p_{j}p_{j}^{\prime}$, with the longest border $b_{j}$, and each prefix of $r_{j}$ with length at least $|p_{j}|$ is closed. Since $|p_{j}^{\prime}|=|b_{j}|-(|b_{j-1}|+1)$, we have $|r_{j}|=|p_{j}|+|p_{j}^{\prime}|=|p_{j}|+|b_{j}|-(|b_{j-1}|+1)$. Solving for $|p_{j}|$, we obtain the desired formula: $|p_{j}|=|r_{j}|-|b_{j}|+|b_{j-1}|+1$. ∎
@@
Theorem 3.1 follows directly from Equation (1), Lemma 2 and 3.
@@
$\mathcal{C}(w)=\{(i,|p_{j}|,|r_{j}|)\mid 1\leq i\leq n,\ 1\leq j\leq K_{i}\},$ (3)
where $K_{i}$ is the number of maximal right-closed substrings starting at position $i$, and for each $j$-th such substring $r_{j}$, starting at index $i$, $p_{j}$ is the shortest closed prefix of $r_{j}$ such that $p_{j}=r_{j}$ or $p_{j}\in E_{r_{j}}$.
@@
$\text{Closed}(w)=\bigcup_{(i,|p_{j}|,|r_{j}|)\in\mathcal{C}(w)}\left\{w[i\ldots i+\ell-1]\mid\ell\in[|p_{j}|,|r_{j}|]\right\}.$ (4)